OpenStudy (just_one_last_goodbye):

which statement is true about the sum √50 + √18? A. It is rational and equal to 2 B. It is rational and equal to 8 C. It is irrational and equal to 2√2 D. It is irrational and equal to 8√2

OpenStudy (mle):

try simplifying the two given roots first sqrt(50) = sqrt(25 * 2) because 25 times 2 = 50 then the sqrt(25) = ? and attach the sqrt(2) onto it

OpenStudy (zzr0ck3r):

\(\sqrt{50}+\sqrt{18}=\sqrt{25*2}+\sqrt{9*2}=\sqrt{25}\sqrt{2}+\sqrt{9}\sqrt{2}\)

OpenStudy (zzr0ck3r):

see if that helps

OpenStudy (just_one_last_goodbye):

I got √2 * 8